The following CVP income statements are available for Blanc Company and Noir Company Blanc Company Noir Company Sales Variable costs Contribution margin Fixed costs Net income $494,000 286,520 207,480 165,000 $42,480 $494,000 192,660 301,340 258,860 $42,480 Calculate Contribution margin ratio. (Round answers to 2 decimal places, e.g. 0.32.) Contribution Margin Ratio Blanc Company Noir Company LINK TO TEXT VIDEO: SIMILAR PROBLEM Compute break-even point in dollars for each company. (Round answers to 0 decimal places, e.g. 1,225.) Break-even Point Blanc Company s Noir Company
Answer #1

Answers

  • All working forms part of the answer
  • Answers are being provided sequentially as asked in picture:

Working

Blanc Company

Noir Company

A

Contribution margin

$          207,480.00

$        301,340.00

B

Sales

$          494,000.00

$        494,000.00

C = A/B

Contribution margin ratio = ANSWER

0.42

0.61

Working

Blanc Company

Noir Company

A

Fixed Costs

$                165,000

$              258,860

B

Contribution margin ratio [calculated above]

0.42

0.61

C = A/B

Break Even point in Dollars = ANSWER

$                392,857

$              424,361

Working

Blanc Company

Noir Company

A

Sales

$                494,000

$              494,000

B

Break Even Sales

$                392,857

$              424,361

C = A - B

Margin of Safety in Dollars

$                101,143

$                 69,639

D = (C/A)

Margin of Safety ratio = ANSWER

                          0.20

                         0.14

Working

Blanc Company

Noir Company

A

Contribution margin

$                207,480

$              301,340

B

Net Income

$                  42,480

$                 42,480

C = A/B

Degree of Operating Leverage = ANSWER

4.9

7.1

---CVP Statement if sales increases by 20%

Working

Blanc Company

Noir Company

A

Sales revenue

$                592,800

$              592,800

B = A - C

Variable cost

$                343,824

$              231,192

C

Contribution margin

$                248,976

$              361,608

D

Fixed Cost

$                165,000

$              258,860

E = C- D

Net Income

$                  83,976

$              102,748

Working for above

Working

Blanc Company

Noir Company

A

Sales revenue

=494000*1.2

=494000*1.2

B = A - C

Variable cost

=592800-248976

=592800-361608

C

Contribution margin

=592800*0.42

=592800*0.61

D

Fixed Cost

165000

258860

E = C- D

Net Income

=248976-165000

=361608-258860

---CVP Statement if Sales decreases by 20%

Working

Blanc Company

Noir Company

A

Sales revenue

$                395,200

$              395,200

B = A - C

Variable cost

$                229,216

$              154,128

C

Contribution margin

$                165,984

$              241,072

D

Fixed Cost

$                165,000

$              258,860

E = C- D

Net Income

$                        984

$              (17,788)

Working for above:

Working

Blanc Company

Noir Company

A

Sales revenue

=494000*0.8

=494000*0.8

B = A - C

Variable cost

=395200-165984

=395200-241072

C

Contribution margin

=395200*0.42

=395200*0.61

D

Fixed Cost

165000

258860

E = C- D

Net Income

=165984-165000

=241072-258860